# InsightEdu: Mobile Discord Bot Management and Analytics for Educators

### v1

@@ Abstract.
Modern educational environments increasingly rely on digital platforms to facilitate interaction between students and educators. Discord has emerged as a popular communication platform in academic settings, offering a combination of messaging and support for chatbot development. However, most existing Discord bots lack specialized educational functionalities and mobile-friendly interfaces, limiting their effectiveness for instructional use. This paper presents InsightEdu, an innovative iOS application that provides a touch-centric interface for managing a custom Discord bot designed for educational contexts. The system enables educators to conduct surveys, collect feedback, and track attendance through an intuitive mobile interface. The architecture combines a SwiftUI-based iOS client application with a Python-based Discord bot server. User evaluation with educators demonstrated significant usability improvements compared to traditional Discord interfaces, with 92% of participants (n = 20) reporting enhanced efficiency in managing educational interactions. This study demonstrates that mobile-first, instructor-friendly design can significantly enhance the utility of existing communication platforms for academic purposes.
@@ 1. Introduction
Interactive learning environments are increasingly essential in computer science education, particularly in large-scale and hybrid courses (interactiveLearning). While several Learning Management Systems (LMS) support content delivery, they often lack immediate in-situ mechanisms for gauging student engagement (moodle). To address this gap, educators have adopted digital communication platforms, most notably Discord, for immediate and contextual interaction due to its flexible channel structure, real-time messaging, and support for custom bots (berrezueta2024interactive). Such platforms can help cultivate transferable soft skills that correlate with early academic success (andree2025softskills). However, Discord’s general-purpose design does not natively support educational management tasks such as attendance tracking, feedback collection, and survey administration (wulanjani2018discord).
Extending Discord with bots is promising. However, current solutions present notable barriers: reliance on cumbersome text-based command interfaces for instructors, limited mobile accessibility, and a lack of specialized educational features (kuhail2023interacting). On the other hand, commercial bot-management dashboards introduce further friction for academic use. Similarly, web-based tools such as Dyno that require OAuth2 flows and browser interfaces and are not optimized for mobile use (dynoWebDashboard2025). At the same time, subscription-oriented services like MEE6 raise concerns about data protection and restrict extensibility through proprietary APIs (mee6Pricing2025). These constraints are problematic as instructors increasingly rely on mobile devices for on-the-go classroom management and real-time analytics.
This paper presents InsightEdu [^1] GitHub repository: https://github.com/TUM-HN/InsightEdu, an open-source iOS application that simplifies Discord-based educational bot management and enables lightweight learning analytics. Building on prior evidence from the TUMDiscordBot research prototype—which demonstrated how slash commands can support attendance tracking, tutor-session feedback, and multi-step surveys in introductory programming courses (berrezueta2024interactive), InsightEdu removes the barrier of command-line workflows and JSON-encoded configurations by providing an intuitive, touch-centric interface. Its architecture decouples bot logic from presentation via a RESTful API gateway, TUMDiscordBot-API, preserving proven pedagogical functionality while enabling secure, mobile-optimized control.
@@
  • Survey administration: Create and distribute simple and multi-step surveys.
The remainder of this paper is organized as follows. Section 2 surveys educational chatbots and mobile EdTech design. Section 3 details the InsightEdu system architecture and implementation, covering the SwiftUI iOS client, the py-cord Discord bot, the Flask REST API, the communication protocol, security and permissions, repository structure, and deployment topology. Section 4 describes the study design, tasks, and metrics used to assess technical performance and usability. Section 5 reports quantitative efficiency, error rates, and latency. Section 6 reflects on implications for educational technology, architectural lessons, limitations, and future research directions. Finally, Section 7 summarizes contributions and outlines avenues for extending InsightEdu.
@@ 2. Related Work
The integration of chatbots in educational environments has gained significant traction as institutions seek to enhance student engagement and collect actionable learning analytics.
2.1. Educational Chatbots and Learning Analytics
Kuhail et al. (kuhail2023interacting) conducted a systematic review of educational chatbot interactions, identifying key success factors including intuitive interfaces, immediate feedback mechanisms, and seamless integration with existing learning workflows. Their analysis revealed that technical complexity is often the primary barrier to chatbot adoption among educators, particularly when deployment requires command-line expertise or complex configuration procedures.
Berrezueta-Guzman et al. (berrezueta2024interactive) demonstrated the pedagogical effectiveness of Discord-based chatbots in computer science education, showing that interactive learning supported by bot-mediated feedback collection improved student engagement and instructor responsiveness to learning difficulties. This study revealed the potential for real-time analytics to inform pedagogical decisions, yet it highlighted the persistent challenge of making such tools accessible to instructors who lack technical backgrounds.
On the other hand, traditional LMS deployments incur substantial costs, with Moodle Cloud standard plans ranging from approximately USD 3,000 to 21,360 annually (moodleCloudPricing2025), excluding maintenance and technical support. Budget constraints in developing regions hinder many institutions from adopting such advanced learning analytics tools. However, Discord’s free hosting model and minimal hardware requirements for bot deployment present a compelling alternative for cost-sensitive educational contexts.
2.2. Mobile Application Development for Education
The evolution of mobile frameworks has revolutionized the development of educational technology, enabling rapid prototyping and cross-platform deployment through declarative UI paradigms. Apple’s SwiftUI represents a significant advancement in iOS development, offering live previews, built-in accessibility support, and seamless integration with platform-specific features (appleSwiftUIDoc2025). Unlike cross-platform frameworks that introduce abstraction layers, SwiftUI compiles to native controls, preserving platform aesthetics and performance characteristics essential for professional educational applications (appleSwiftUITut2025).
The introduction of SwiftData (wwdc2023SwiftData) further streamlined iOS data persistence, providing code-first model definitions that integrate directly with SwiftUI’s reactive data flow. This advancement eliminates much of the boilerplate code historically required by Core Data [^2] Apple’s Core Data helps apps keep and organize their data, supporting local modeling, storage, and search on iOS (appleCoreData)., enabling developers to focus on user experience rather than infrastructure. SwiftData’s automatic schema migration and Combine [^3] Combine is Apple’s reactive framework with a declarative Swift API for handling values over time—like async events and data streams (appleCombine). integration offer advantages over traditional persistence approaches for educational applications requiring offline capability and secure local storage.
Mobile application architecture patterns have evolved to support modern UI frameworks, enabling seamless integration with these frameworks. The Model-View-ViewModel (MVVM) pattern is the preferred approach for SwiftUI applications (anderson2012model; appeloper2022stop), offering an optimal balance between testability and code maintainability. While alternatives such as Model-View-Controller (MVC) (garcia2023mvc) or VIPER (vennaro2023viper) exist, MVVM’s reactive binding model aligns naturally with SwiftUI’s declarative syntax, reducing impedance mismatch between architectural patterns and framework capabilities (aljamea2018mmvmi).
@@ 3.1. Overview and Requirements
The system comprises three core components: (1) a SwiftUI-based iOS client, (2) a Flask REST API, and (3) a Python Discord bot built with py-cord. Figure 1 provides a high-level view of these components and their interactions.
Figure 1. High-level system architecture showing the iOS client application, Flask REST API, Discord bot, Discord Gateway, and shared storage.
Beyond documenting instructors’ command-line pain points—such as remembering slash-command syntax and hand-editing JSON—we also identified a lack of real-time feedback during lectures and a reliance on external tools for post-hoc data processing. We addressed these by refining the functional (FR) and nonfunctional (NFR) requirements in workshops with tutor cohorts and aligning them with capabilities already validated by TUMDiscordBot:
@@
  • FR-2: Create simple one-question surveys with button/emoji reactions.
  • FR-3: Create complex multi-question surveys with free-text responses.
@@
  • NFR-2: The UI conforms to Apple’s Human Interface Guidelines (appleHIG2024);
@@
Our reference setup runs on a Raspberry Pi (4GB RAM) kept on campus. It keeps costs low while handling typical classes of 100–150 students per session.
3.2. iOS Client Application (SwiftUI, Model-View-ViewModel/MVVM)
The iOS client is implemented in SwiftUI and follows a pragmatic, Model-View-ViewModel (MVVM) pattern (appeloper2022stop). Each view manages local state (@State), orchestrates network operations via a centralized APIClient (wrapping URLSession with async/await), and persists domain models using SwiftData’s @Model macro. The app targets iOS 18.2+ and supports iPhone/iPad with adaptive layouts and deep linking to specific flows (ios182Release). This reduces boilerplate while separating concerns across presentation, domain, and data layers, as illustrated in Figure 2.
Figure 2. Internal app layers: presentation (views, charts), domain models, and data (APIClient, persistence).
3.2.1. User Interface Design
The UI is touch-centric and adheres to Apple’s Human Interface Guidelines (HIG) (appleHIG2024):
@@ 3.2.1. User Interface Design
  • Data visualization: Charts for feedback/survey summaries (Apple Charts (appleCharts2024)).
Figure 3 presents the main screen of the InsightEdu iOS application, which provides educators with direct access to educational bot management features.
Figure 3. Main screen of the InsightEdu iOS application with access to educational features.
@@
  • Title: DiscordBot Manager (centered at the top).
  • Status Information: Displays the current server status (“Server Online”), the server URL, and the status check time.
  • Settings button: Located in the top-right corner.
@@
  • Server Stats: Three small panels showing the number of users Online, Offline, and Total.
  • Action Buttons:
    • Delete Bot (red button on the right).
    • Start Bot (green button on the right).
UI Elements - Command Categories
The commands are grouped into functional blocks, each presented as a tappable list item with icons and labels:
  • Simple Commands:
    • Ping: Check bot response time.
  • Member Commands:
    • Send Message: Send a personalized greeting.
    • Give Role: Assign a role to a member.
  • Channel Commands:
    • Clear Messages: Delete messages from a channel.
  • Group Commands:
    • Attendance: Track attendance in a channel.
    • Instructor Feedback: Collect session feedback.
  • Survey Commands:
    • Simple Survey: Create a single-question survey.
    • Complex Survey: Create a multi-question survey.
The UI is hierarchical and task-oriented, with large, clearly labeled buttons for quick navigation. A short description accompanies each feature to reduce cognitive load.
3.2.2. App Architecture, Navigation, and Data Management
Navigation utilizes a hierarchical flow from a single landing screen (illustrated in Figure 4); touch flows are optimized for frequent actions (see Figure 5). Domain objects (e.g., Bot, GroupModel, SurveyFile) are persisted with SwiftData, enabling offline preparation and local history. Sensitive secrets (API keys, tokens) leverage SwiftData/keychain integration for General Data Protection Regulation (GDPR) aligned storage (gdpr2016).
Figure 4. Navigation hierarchy from the landing screen to feature flows.
Figure 5. Touch-centric flow optimized for minimal taps for common teaching tasks.
Network Data Transfer Objects (DTOs) mirror server JSON schemas via Swift’s Codable [^4] Codable is a Swift protocol that lets you easily turn types into JSON and back.; the APIClient exposes typed methods (e.g., createComplexSurvey(...)), performs non-blocking requests with async/await, and applies retry with exponential backoff to handle transient errors. Charts are updated with new data from the server to present real-time lecture feedback.
@@ 3.3. Server Architecture and REST API
The server uses Flask (flask2024) and a dual-threaded design: a REST thread handles HTTP requests and authentication, while a py-cord event loop maintains a persistent Discord Gateway connection. This preserves responsiveness during long-running Discord operations without IPC complexity. Figure 6 depicts the server-side components and data flows.
Figure 6. Server-side architecture with API service, bot worker, and data stores behind a reverse proxy.
Flask blueprints organize endpoints by domain, following the Level 2 REST conventions. The client communicates via a small, versioned set of JSON endpoints; requests use short-lived bearer tokens or API keys (see Security), with RFC 7807-style error payloads. A representative survey-creation interaction is shown in Figure 7.
Figure 7. Representative REST interaction between the iOS client and the server for survey creation and result retrieval.
Error handling standardizes JSON responses and status codes (e.g., 400 for validation errors, 403 for authentication errors, 500 for server errors). Audit logging uses structured JSON with daily rotation; CSV exports (data/attendance/*.csv, data/surveys/*.csv) support institutional analytics and manual inspection.
3.4. Communication Protocol and Endpoint Example
A typical “Start Attendance” flow:
  1. (1)
    Instructor taps Start Attendance in AttendanceCommandView.
  2. (2)
    App issues POST /api/attendance?
    code=1423&group=g1&status=start
  3. (3)
    Flask validates credentials and schedules the coroutine on the bot loop.
  4. (4)
    The bot posts an attendance prompt; students reply with the code.
  5. (5)
    Flask returns a successful JSON response; the app displays a confirmation toast.
Example response payload:
  "status": "success",
  "message": "Attendance command executed..."
3.5. Security and Permissions
Transport & Perimeter.
TLS termination, CORS, security headers, and rate limiting are handled by an NGINX reverse proxy (nginxDocs2024). Let’s Encrypt certificates automate certificate management (letsencrypt2025). Security headers follow OWASP ASVS guidance (owaspASVS2023).
API Authentication.
A flat API-key scheme is used for a small, trusted instructor cohort: short opaque strings stored in an encrypted .secrets.json and validated by decorators on each request. This keeps operations auditable and straightforward while allowing an upgrade path to JWT (jain2025mobileapi) (scoped access, expiration) without architectural changes.
Discord Permissions.
The bot leverages Discord’s granular permissions but contracts administrative capabilities around a single Admin role. Each REST-triggered action executes under a mock context carrying that role, preserving native permission checks.
Privacy & Validation.
Inputs are validated to prevent injection; rate limiting constrains abuse. The mobile client minimizes sensitive data by storing configuration locally with SwiftData/keychain integration and gathers user consent where applicable, supporting GDPR-aligned practices.
3.6. Repository Structure
The repository separates the iOS app, API service, and bot worker to streamline onboarding and maintenance, mirroring the layered and deployment split. Figure 8 presents an overview that mirrors the layers discussed earlier and the deployment split shown on the server side.
Figure 8. Repository structure across iOS app, API service, and bot worker.
3.7. Deployment Topology
In the reference deployment, NGINX [^5] NGINX is a high-performance web server and reverse proxy, as well as an IMAP/POP3 proxy server. reverse proxy terminates TLS and forwards HTTP to Flask on a Raspberry Pi. A single Python process runs two threads: (1) Flask REST API and (2) the Discord bot event loop (py-cord 2.5.0). Both share the local filesystem for exports and an encrypted .secrets.json (API keys, role IDs). For heavier loads, the backend can be split into an API service and a bot worker, optionally with a message queue to decouple long-running tasks; services are containerized and managed via systemd (systemdDocs2024) and deployed behind the reverse proxy (Figure 6).
@@ 3.8. Discord Bot Implementation
The Discord bot is implemented with py-cord (pycordDocs2025) and exposes standardized slash commands (discordSlashCommands2021). Command categories include:
  • Attendance: student check-in; instructor start/stop.
  • Feedback: request, submit, and view session feedback.
  • Surveys: create, distribute, and analyze surveys.
These Discord-based interactions are designed to be intuitive for instructors and provide valuable data that instructors can analyze through the InsightEdu. The bot integrates with Discord’s permission and role system, allowing only authorized instructors to run administrative commands. On startup, a shared context is initialized (server URL, API token, selected guild/channels, runtime flags) and made available to cogs [^6] In Discord.py, a cog is a modular class that groups commands and listeners, helping organize bot functionality (discordpy_cogs). and event handlers (Figure 9).
Figure 9. Bot context initialization and shared state for commands and events.
3.8.1. Feature Implementation
@@
Through InsightEdu, instructors can start sessions, monitor real-time check-ins, and review history as illustrated in Figure 10. The workflow proceeds from the session start (via the attendance code for authorization), student check-ins via direct message with this attendance code, and session finalization by the instructor.
Figure 10. Visualization of the attendance at the lecture (the names are blurred for privacy reasons).
@@
InsightEdu collects structured feedback (rating scales, open-text) and visualizes aggregates (see Figure 11).
Figure 11. Visualization of the feedback provided by the students immediately after a tutor session.
@@
For more detailed feedback, instructors can create complex surveys that initiate in the channel (see Figure 15) and continue through direct messages after clicking the ”Participate” button, allowing for a sequence of questions with either percentage or verbal difficulty level response types (see Figure 16).
Figure 15. Complex survey initiation in a Discord channel, where students click ’Participate’ to begin the survey process.
Figure 16. Complex survey questions delivered through Discord direct messages, allowing for sequential responses.
The results of these complex surveys are also presented to the instructors in similar charts, as shown in Figure 16.
@@ 4. Evaluation
The evaluation methodology combined quantitative performance measurements with qualitative usability assessment to validate technical functionality and educational utility. Twenty experienced course instructors from various lectures participated in comparative studies that contrasted CLI-based bot management with the mobile interface across representative educational scenarios. Participants possessed varying levels of technical expertise, ranging from computer science faculty to teaching assistants with limited experience using the command line.
Task scenarios encompassed the seven core functional requirements identified during system design:
  • Attendance tracking initiation for tutorial groups (Figure 10),
  • Simple survey creation with emoji reactions,
  • Complex multi-question survey deployment,
  • Tutor session feedback collection (see Figure 11),
  • Content/exercise difficulty perception (see Figure 14)
  • Bot configuration management
  • Historical data analysis through interactive visualizations.
Each scenario was performed using CLI commands and the mobile interface, with completion times, error rates, and subjective satisfaction metrics recorded for comparative analysis.
The evaluation environment replicated typical classroom deployment conditions using campus Wi-Fi networks with measured latency characteristics and a Raspberry Pi 4B (4GB RAM) hosting the bot infrastructure. Network conditions varied throughout the evaluation period, providing a realistic assessment of system performance under variable connectivity scenarios common in educational environments.
Usability evaluation employed heuristic analysis based on Jakob Nielsen’s principles, adapted for mobile educational applications. Assessment criteria included the visibility of system status, the match between the system and the real world, user control and freedom, consistency and standards, error prevention, recognition rather than recall, flexibility and efficiency of use, aesthetic and minimalist design, help and documentation quality, and recovery from error conditions.
Quantitative metrics included task completion time, successful completion rate, number of user errors per task, time to recover from mistakes, and system response latency for individual operations. Qualitative assessment captured subjective satisfaction ratings, perceived cognitive load, preference rankings between interface modalities, and suggestions for functionality improvements or workflow optimization.

### v2

@@ Abstract.
Modern educational environments increasingly rely on digital platforms to facilitate interaction between students and educators. Discord has emerged as a popular communication platform in academic settings, offering a combination of messaging and support for chatbot development. However, most existing Discord bots lack specialized educational functionalities and mobile-friendly interfaces, limiting their effectiveness for instructional use. This paper presents InsightEdu, a mobile-first interaction design and reference architecture for instructor-operated Discord bots. The system enables educators to conduct surveys, collect feedback, and track attendance through an intuitive mobile interface. The architecture combines a SwiftUI-based iOS client application with a Python-based Discord bot server. User evaluations with educators demonstrated significant usability improvements over traditional Discord interfaces, with 92% of participants (n = 20) reporting enhanced efficiency in managing educational interactions. This study demonstrates that mobile-first, instructor-friendly design can significantly enhance the utility of existing communication platforms for academic purposes.
@@ 1. Introduction
Interactive learning environments are increasingly essential in computer science education, particularly in large-scale and hybrid courses (Krusche and Berrezueta-Guzman, 2023). While several Learning Management Systems (LMS) support content delivery, they often lack immediate in-situ mechanisms for gauging student engagement (Gamage et al., 2022). To address this gap, educators have adopted digital communication platforms, most notably Discord, for immediate, context-specific interaction due to its flexible channel structure, real-time messaging, and support for custom bots (Berrezueta-Guzman et al., 2024). Such platforms can help cultivate transferable soft skills that correlate with early academic success (Andree et al., 2025). However, Discord’s general-purpose design does not natively support educational management tasks such as attendance tracking, feedback collection, and survey administration (Wulanjani, 2018).
Extending Discord with bots is promising. However, current solutions present notable barriers: reliance on cumbersome text-based command interfaces for instructors, limited mobile accessibility, and a lack of specialized educational features (Kuhail et al., 2023). On the other hand, commercial bot-management dashboards introduce further friction for academic use. Similarly, web-based tools such as Dyno that require OAuth2 flows and browser interfaces and are not optimized for mobile use (Aquino et al., 2025). At the same time, subscription-oriented services like MEE6 raise concerns about data protection and restrict extensibility through proprietary APIs (MEE6 Community, ). These constraints are problematic as instructors increasingly rely on mobile devices for on-the-go classroom management and real-time analytics.
This paper presents InsightEdu [^1] GitHub repository: https://github.com/TUM-HN/InsightEdu, an open-source iOS application that simplifies Discord-based educational bot management and enables lightweight learning analytics. Building on prior evidence from the TUMDiscordBot research prototype—which demonstrated how slash commands can support attendance tracking, tutor-session feedback, and multi-step surveys in introductory programming courses (Berrezueta-Guzman et al., 2024), InsightEdu removes the barrier of command-line workflows and JSON-encoded configurations by providing an intuitive and touch-centric interface. Its architecture decouples bot logic from presentation via a RESTful API gateway, TUMDiscordBot-API, preserving proven pedagogical functionality while enabling secure, mobile-optimized control.
@@
  • Survey administration: Create and distribute simple and complex surveys.
The remainder of this paper is organized as follows. Section 2 surveys educational chatbots and mobile EdTech design. Section 3 details the InsightEdu system architecture and implementation. Section 4 describes the study design, tasks, and metrics used to assess technical performance and usability. Section 5 reports quantitative efficiency, error rates, and latency. Section 6 reflects on implications for educational technology, architectural lessons, limitations, and future research directions. Finally, Section 7 concludes the paper.
@@ 2. Related Work
The integration of educational chatbots has gained traction as a tool to enhance student engagement and collect learning analytics. Kuhail et al. (Kuhail et al., 2023) identified that technical complexity remains the primary barrier to chatbot adoption among educators. While Discord-based bots have proven effective in improving student-instructor interaction in computer science courses, existing solutions often require command-line expertise for deployment and configuration (Berrezueta-Guzman et al., 2024). Furthermore, traditional Learning Management Systems (LMS) like Moodle Cloud incur substantial annual costs (up to USD 21,360) (MoodleCloud, 2025). Discord’s free hosting model provides a cost-effective alternative for budget-constrained institutions.
Modern declarative UI frameworks like SwiftUI have revolutionized educational technology by enabling rapid prototyping with native performance and built-in accessibility (Apple Inc., 2025b, c). SwiftData further streamlines this by providing code-first data persistence that integrates with SwiftUI’s reactive data flow, replacing the high boilerplate requirements of legacy frameworks like Core Data (Apple Inc., 2023, 2025a).
Model-View-ViewModel (MVVM) is the preferred architectural pattern for SwiftUI applications (Magics-Verkman et al., 2023). Unlike MVC or VIPER, MVVM’s reactive binding model aligns naturally with declarative syntax, reducing the impedance mismatch between logic and presentation layers (Anderson, 2012; Aljamea and Alkandari, 2018).
@@ 3.1. Overview and Requirements
The system comprises three core components: (1) a SwiftUI-based iOS client, (2) a Flask REST API, and (3) a Python Discord bot built with py-cord.
Beyond documenting instructors’ command-line pain points, such as remembering slash, command syntax, and hand-editing JSON, we also identified a lack of real-time feedback during lectures and a reliance on external tools for post-hoc data processing. We addressed these by refining the functional (FR) and nonfunctional (NFR) requirements in workshops with tutor cohorts and aligning them with capabilities already validated by TUMDiscordBot:
@@
  • FR-2: Simple surveys with button/emoji reactions.
  • FR-3: Complex multi-question surveys with free-text responses.
@@
  • NFR-2: GUI follows Apple’s Human Interface Guidelines;
@@
Our bot setup runs on a Raspberry Pi (4GB RAM) kept on campus. It keeps costs low while handling typical classes of 100–150 students per session.
3.2. User Interface Design
The UI is touch-centric and adheres to Apple’s Human Interface Guidelines (HIG) (Apple Inc., 2024a):
@@ 3.2. User Interface Design
  • Data visualization: Charts for feedback/survey summaries (Apple Charts (Apple Inc., 2024b)).
Figure 1 presents the main screen of the InsightEdu iOS application, which provides educators with direct access to educational Discord bot management features. Among its design we have:
@@
  • Title, Status Information, and Settings
@@
  • Server Stats, Action Buttons (Delete Bot, Start Bot).
UI Elements - Command Categories The commands are grouped into functional blocks:
  • Simple Commands: Ping:
  • Member Commands: Send Message and Give Role
  • Channel Commands: Clear Messages
  • Group Commands: Attendance and Instructor Feedback
  • Survey Commands: Simple Survey and Complex Survey
Figure 1. Main screen of the InsightEdu iOS application.
The UI is hierarchical and task-oriented, with large, clearly labeled buttons for quick navigation. Additionally, a short description accompanies each feature to reduce cognitive load.
@@ 3.3. Server Architecture and REST API
The server uses Flask (Pallets Project, 2024) and a dual-threaded design, a REST thread handles HTTP requests and authentication, while a py-cord event loop maintains a persistent Discord Gateway connection.
Flask blueprints organize endpoints by domain, following the Level 2 REST conventions. The client communicates via a small, versioned set of JSON endpoints; requests use short-lived bearer tokens or API keys, with RFC 7807-style error payloads.
Error handling standardizes JSON responses and status codes. Audit logging uses structured JSON with daily rotation; CSV exports (data/attendance/*.csv, data/surveys/*.csv) support institutional analytics and manual inspection.
@@ 3.4. Discord Bot Implementation
The Discord bot is implemented with py-cord (Pycord Development Team, 2025) and exposes standardized slash commands (Discord Inc., 2021). Command categories include:
@@
Through InsightEdu, instructors can start sessions, monitor real-time check-ins, and review history. The workflow proceeds from the session start (via the attendance code for authorization), student check-ins via direct message with this attendance code, and session finalization by the instructor.
@@
InsightEdu collects structured feedback (rating scales, open-text) and visualizes aggregates.
@@
For more detailed feedback, instructors can create complex surveys that begin in the channel and continue via direct messages after clicking the ”Participate” button, allowing a sequence of questions (see Figure 5). The results of these complex surveys are also presented to the instructors in similar charts, as shown in Figure 4.
Figure 5. Complex survey in a Discord channel.
@@ 4. Evaluation
The evaluation validated InsightEdu through a comparative study contrasting traditional CLI-based bot management with the mobile interface. Twenty experienced instructors with varying technical backgrounds performed six core tasks: attendance tracking, complex and straightforward survey deployment, feedback collection, configuration management, and historical data analysis.
Tests were conducted using a Raspberry Pi 4B over campus Wi-Fi to simulate realistic classroom environments. Usability was assessed using Jakob Nielsen’s heuristics, while cognitive load was measured via the NASA-TLX index. Quantitative metrics included task completion time, error rates, and system response latency. Qualitative assessment captured subjective satisfaction and interface preferences through structured interviews and rankings.